# The theory of implicit operations

### v1

@@ 7. Tangible epimorphism surjectivity
Now, we prove that $\mathsf{RCDL}$ has the strong epimorphism surjectivity property. As $\mathsf{RCDL}$ has a lattice reduct, it possesses a majority term. Therefore, we can apply 7.3, obtaining that $\mathsf{RCDL}$ has the strong epimorphism surjectivity property if and only if for each finitely generated $\bm{A}\leqslant\bm{B}\times\bm{C}$ with $\bm{B},\bm{C}\in\mathsf{RCDL}_{\textsc{fsi}}$ finitely generated we have $\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B}\times\bm{C})=A$. Together with 7.6 the latter specializes to the following: for each $\bm{A}\leqslant\bm{D}_{2}\times\bm{D}_{2}$ we have $\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{D}_{2}\times\bm{D}_{2})=A$.
@@
As $\varphi$ defines an implicit operation of $\mathsf{K}$, it also defines an implicit operation $g$ of $\mathbb{Q}(\mathsf{K})$ by 3.11. As both $f$ and $g$ are defined by $\varphi$ and $\bm{D}\in\mathsf{K}\subseteq\mathbb{Q}(\mathsf{K})$, we have $f^{\bm{D}}=g^{\bm{D}}$. We will make use of this observation without further notice.
@@
Define $\bm{B}=\bm{T}_{\mathbb{Q}(\mathsf{K})}(\varphi)$, $a_{1}=x_{1}/\theta(\varphi),\dots,a_{n}=x_{n}/\theta(\varphi)$, and $\bm{A}=\mathsf{Sg}^{\bm{B}}(a_{1},\dots,a_{n})$ (for the definition of $\bm{T}_{\mathbb{Q}(\mathsf{K})}(\varphi)$ see 4.12). Clearly, $\bm{A}$ is a finitely generated subalgebra of $\bm{B}$. Moreover, from 4.13(i) it follows that $\bm{B}$ is finitely presented in $\mathbb{Q}(\mathsf{K})$ and that
@@
Now, recall from (23) that $\langle c_{1},\dots,c_{n}\rangle\in\mathsf{dom}(f^{\bm{D}})=\mathsf{dom}(g^{\bm{D}})$. Together with $\bm{D}\in\mathsf{K}\subseteq\mathbb{Q}(\mathsf{K})$ and 4.13(ii), this guarantees the existence of a homomorphism $h\colon\bm{B}\to\bm{D}$ such that
@@
Now, we proceed to prove 7.3. We will make use of the next concept from [29, Def. 4.4].
@@
As $\bm{A}\leqslant\bm{B}\in\mathsf{K}$ and $b\in\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})-A\subseteq B-A$, we can apply 7.10, obtaining $\phi\in\mathsf{Con}_{\mathsf{K}}(\bm{B})$ satisfying the following requirements: $b/\phi\in B/\phi-A/\phi$ and for each $\theta\in\mathsf{Con}_{\mathsf{K}}(\bm{B}/\phi)-\{\textup{id}_{B/\phi}\}$ there exists $a\in A$ such that $\langle a/\phi,b/\phi\rangle\in\theta$.
Clearly, $\bm{A}/\phi\leqslant\bm{B}/\phi$ is a proper subalgebra. Moreover, $\bm{A}/\phi$ and $\bm{B}/\phi$ are finitely generated members of $\mathsf{K}$ because so are $\bm{A}$ and $\bm{B}$ by assumption and $\phi\in\mathsf{Con}_{\mathsf{K}}(\bm{B})$. 4.6(ii) implies that $b/\phi\in\mathsf{d}_{\mathsf{K}}(\bm{A}/\phi,\bm{B}/\phi)$. As $b/\phi\notin A/\phi$, we obtain $b/\phi\in\mathsf{d}_{\mathsf{K}}(\bm{A}/\phi,\bm{B}/\phi)-A/\phi$. Therefore, we may assume that $\phi=\textup{id}_{B}$ (otherwise we replace $\bm{A}$ and $\bm{B}$ by $\bm{A}/\phi$ and $\bm{B}/\phi$, respectively).
@@
Let $\theta_{1},\dots,\theta_{n}\in\mathsf{Con}_{\mathsf{K}}(\bm{B})$ be such that $\text{id}_{B}=\theta_{1}\cap\dots\cap\theta_{n}$. Let also $\phi_{i}=\theta_{1}\cap\dots\cap\theta_{i-1}\cap\theta_{i+1}\dots\cap\theta_{n}$ for each $i\leqslant n$. We will show that $\phi_{i}=\text{id}_{B}$ for some $i\leqslant n$. Suppose the contrary, with a view to contradiction. 7.11 yields $a_{1},\dots,a_{n}\in A$ such that $\langle a_{i},b\rangle\in\phi_{i}$ for every $i\leqslant n$. By assumption $\mathsf{K}$ has a near unanimity term $t(x_{1},\dots,x_{n})$. We will prove that
@@
In view of Claim 7.12 and Proposition 7.9, there exist $\theta_{1},\dots,\theta_{n-1}\in\mathsf{Irr}_{\mathsf{K}}(\bm{B})$ such that $\text{id}_{B}=\theta_{1}\cap\dots\cap\theta_{n-1}$. Therefore, we can apply Proposition 2.7 obtaining a subdirect embedding $h\colon\bm{B}\to\bm{B}/\theta_{1}\times\dots\times\bm{B}/\theta_{n-1}$. Let $\bm{B}_{i}=\bm{B}/\theta_{i}$ for each $i\leqslant n-1$. By replacing $\bm{A}$ and $\bm{B}$ by their isomorphic images $h[\bm{A}]$ and $h[\bm{B}]$, respectively, we may assume that $\bm{A}\leqslant\bm{B}\leqslant\bm{B}_{1}\times\dots\times\bm{B}_{n-1}$. Notice that each $\bm{B}_{i}=\bm{B}/\theta$ is finitely generated because so is $\bm{B}$. Furthermore, from Proposition 2.10 and $\theta_{i}\in\mathsf{Irr}_{\mathsf{K}}(\bm{B})$ it follows that $\bm{B}_{i}\in\mathsf{K}_{\textsc{rfsi}}$. Lastly, as $b\in\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})-A$ and $\bm{B}\leqslant\bm{B}_{1}\times\dots\times\bm{B}_{n-1}$, 4.6(ii) allows us to conclude that $b\in\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B}_{1}\times\dots\times\bm{B}_{n-1})-A$. Hence, $\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B}_{1}\times\dots\times\bm{B}_{n-1})\neq A$. ∎
The literature on epimorphisms contains two variants of Theorem 7.3 in which the class $\mathsf{K}$ is required to be an arithmetical variety with the property that the class of its finitely subdirectly irreducible members is closed under ultraproducts and nontrivial subalgebras [23, Thm. 6.8] or only a congruence permutable variety [29, Thm. 5.3]. The first variant deals with the demand that all $\mathsf{K}$-epimorphisms be surjective, while the second with the weaker demand that all $\mathsf{K}$-epimorphisms between finitely generated algebras be surjective called the weak epimorphism surjectivity property. In both cases, the conclusion is that failures of the relevant property are witnessed by counterexamples of the form $\bm{A}\leqslant\bm{B}$ where $\bm{B}$ is a finitely subdirectly irreducible member of $\mathsf{K}$. The possibility of obtaining similar results for the strong epimorphism surjectivity property is prevented by the following example. However, we will show in 7.16 that, under the amalgamation property, the above mentioned result for congruence permutable varieties becomes available in the context of the strong epimorphism surjectivity property as well.
@@
The next result is well known (see, e.g., [14, Thm. 1.3]). We provide a novel and short proof using the characterization of dominions in the presence of the amalgamation property established in 4.14.
@@
The implication from left to right is straightforward. So, let us assume that $\mathsf{K}$ has the weak epimorphism surjectivity property. We will show that $\mathsf{K}$ has the strong epimorphism surjectivity property using Proposition 6.6. To this end, consider $\bm{A}\leqslant\bm{B}\in\mathsf{K}$ and $b\in\mathsf{dom}_{\mathsf{K}}(\bm{A},\bm{B})$. Since $\mathsf{K}$ has the amalgamation property, by 4.14 there exist $f\in\mathsf{imp}_{eq}(\mathsf{K})$ and $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{B}})\cap A^{n}$ such that $f^{\bm{B}}(a_{1},\dots,a_{n})=b$. Let $\bm{A}^{\prime}=\mathsf{Sg}^{\bm{A}}(a_{1},\dots,a_{n})$ and $\bm{B}^{\prime}=\mathsf{Sg}^{\bm{B}}(a_{1},\dots,a_{n},b)$. Since $f$ is defined by a conjunction of equations, $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{B}^{\prime}})\cap(A^{\prime})^{n}$ and $f^{\bm{B}^{\prime}}(a_{1},\dots,a_{n})=b$. So, 4.14 implies that $b\in\mathsf{d}_{\mathsf{K}}(\bm{A}^{\prime},\bm{B}^{\prime})$. As $\bm{B}^{\prime}=\mathsf{Sg}^{\bm{B}}(a_{1},\dots,a_{n},b)$ and $a_{1},\dots,a_{n},b\in\mathsf{d}_{\mathsf{K}}(\bm{A}^{\prime},\bm{B}^{\prime})$, we obtain $\mathsf{d}_{\mathsf{K}}(\bm{A}^{\prime},\bm{B}^{\prime})=B^{\prime}$. Therefore, the inclusion map $\bm{A}^{\prime}\to\bm{B}^{\prime}$ is an epimorphism. Since $\bm{A}^{\prime},\bm{B}^{\prime}$ are finitely generated members of $\mathsf{K}$ and $\mathsf{K}$ has the weak epimorphism surjectivity property, it follows that $\bm{A}^{\prime}=\bm{B}^{\prime}$, and hence $b\in B^{\prime}=A^{\prime}\subseteq A$. We have shown that $\mathsf{dom}_{\mathsf{K}}(\bm{A},\bm{B})=A$. Thus, $\mathsf{K}$ has the strong epimorphism surjectivity property. ∎
@@
The implication from left to right is straightforward. On the other hand, if every finitely generated $\bm{B}\in\mathsf{K}_{\textsc{fsi}}$ lacks proper $\mathsf{K}$-epic subalgebras, then 7.15 guarantees that $\mathsf{K}$ has the weak epimorphism surjectivity property, which by 7.14 implies that $\mathsf{K}$ has the strong epimorphism surjectivity property as well. ∎
@@
Assume that $\mathsf{K}=\mathbb{V}(\mathsf{K}_{1}\cup\dots\cup\mathsf{K}_{n})$, where each $\mathsf{K}_{i}$ is a variety with the weak epimorphism surjectivity property. Suppose, with a view to contradiction, that $\mathsf{K}$ lacks this property. By 7.15 this implies that there exists a finitely generated $\bm{B}\in\mathsf{K}_{\textsc{fsi}}$ with a finitely generated subalgebra $\bm{A}\leqslant\bm{B}$ that is proper and $\mathsf{K}$-epic. Applying 2.12, we obtain that $\bm{B}\in\mathbb{H}\mathbb{S}\mathbb{P}_{\!\textsc{{u}}}(\mathsf{K}_{1}\cup\dots\cup\mathsf{K}_{n})$. By [11, Thm. 5.6] we have $\mathbb{P}_{\!\textsc{{u}}}(\mathsf{K}_{1}\cup\dots\cup\mathsf{K}_{n})=\mathbb{P}_{\!\textsc{{u}}}(\mathsf{K}_{1})\cup\dots\cup\mathbb{P}_{\!\textsc{{u}}}(\mathsf{K}_{n})$. Therefore, $\bm{B}\in\mathbb{H}\mathbb{S}\mathbb{P}_{\!\textsc{{u}}}(\mathsf{K}_{1})\cup\dots\cup\mathbb{H}\mathbb{S}\mathbb{P}_{\!\textsc{{u}}}(\mathsf{K}_{n})\subseteq\mathsf{K}_{1}\cup\dots\cup\mathsf{K}_{n}$. Then $\bm{B}\in\mathsf{K}_{i}$ for some $i\leqslant n$. But $\mathsf{K}_{i}$ has the weak epimorphism surjectivity property by assumption, whence $\bm{A}\leqslant\bm{B}$ cannot be a $\mathsf{K}_{i}$-epic subalgebra. As $\mathsf{K}_{i}\subseteq\mathsf{K}$, in particular it follows that $\bm{A}\leqslant\bm{B}$ cannot be a $\mathsf{K}$-epic subalgebra either. But this contradicts the assumption and thus completes the proof. ∎

### v2

@@ 7. Tangible epimorphism surjectivity
Now, we prove that $\mathsf{RCDL}$ has the strong epimorphism surjectivity property. As $\mathsf{RCDL}$ has a lattice reduct, it possesses a majority term. Therefore, we can apply Theorem 7.3, obtaining that $\mathsf{RCDL}$ has the strong epimorphism surjectivity property if and only if for each finitely generated $\bm{A}\leqslant\bm{B}\times\bm{C}$ with $\bm{B},\bm{C}\in\mathsf{RCDL}_{\textsc{fsi}}$ finitely generated we have $\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B}\times\bm{C})=A$. Together with Claim 7.6 the latter specializes to the following: for each $\bm{A}\leqslant\bm{D}_{2}\times\bm{D}_{2}$ we have $\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{D}_{2}\times\bm{D}_{2})=A$.
@@
As $\varphi$ defines an implicit operation of $\mathsf{K}$, it also defines an implicit operation $g$ of $\mathbb{Q}(\mathsf{K})$ by Corollary 3.11. As both $f$ and $g$ are defined by $\varphi$ and $\bm{D}\in\mathsf{K}\subseteq\mathbb{Q}(\mathsf{K})$, we have $f^{\bm{D}}=g^{\bm{D}}$. We will make use of this observation without further notice.
@@
Define $\bm{B}=\bm{T}_{\mathbb{Q}(\mathsf{K})}(\varphi)$, $a_{1}=x_{1}/\theta(\varphi),\dots,a_{n}=x_{n}/\theta(\varphi)$, and $\bm{A}=\mathsf{Sg}^{\bm{B}}(a_{1},\dots,a_{n})$ (for the definition of $\bm{T}_{\mathbb{Q}(\mathsf{K})}(\varphi)$ see Definition 4.12). Clearly, $\bm{A}$ is a finitely generated subalgebra of $\bm{B}$. Moreover, from Proposition 4.13(i) it follows that $\bm{B}$ is finitely presented in $\mathbb{Q}(\mathsf{K})$ and that
@@
Now, recall from (23) that $\langle c_{1},\dots,c_{n}\rangle\in\mathsf{dom}(f^{\bm{D}})=\mathsf{dom}(g^{\bm{D}})$. Together with $\bm{D}\in\mathsf{K}\subseteq\mathbb{Q}(\mathsf{K})$ and Proposition 4.13(ii), this guarantees the existence of a homomorphism $h\colon\bm{B}\to\bm{D}$ such that
@@
Now, we proceed to prove Theorem 7.3. We will make use of the next concept from [29, Def. 4.4].
@@
As $\bm{A}\leqslant\bm{B}\in\mathsf{K}$ and $b\in\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})-A\subseteq B-A$, we can apply Proposition 7.10, obtaining $\phi\in\mathsf{Con}_{\mathsf{K}}(\bm{B})$ satisfying the following requirements: $b/\phi\in B/\phi-A/\phi$ and for each $\theta\in\mathsf{Con}_{\mathsf{K}}(\bm{B}/\phi)-\{\textup{id}_{B/\phi}\}$ there exists $a\in A$ such that $\langle a/\phi,b/\phi\rangle\in\theta$.
Clearly, $\bm{A}/\phi\leqslant\bm{B}/\phi$ is a proper subalgebra. Moreover, $\bm{A}/\phi$ and $\bm{B}/\phi$ are finitely generated members of $\mathsf{K}$ because so are $\bm{A}$ and $\bm{B}$ by assumption and $\phi\in\mathsf{Con}_{\mathsf{K}}(\bm{B})$. Corollary 4.6(ii) implies that $b/\phi\in\mathsf{d}_{\mathsf{K}}(\bm{A}/\phi,\bm{B}/\phi)$. As $b/\phi\notin A/\phi$, we obtain $b/\phi\in\mathsf{d}_{\mathsf{K}}(\bm{A}/\phi,\bm{B}/\phi)-A/\phi$. Therefore, we may assume that $\phi=\textup{id}_{B}$ (otherwise we replace $\bm{A}$ and $\bm{B}$ by $\bm{A}/\phi$ and $\bm{B}/\phi$, respectively).
@@
Let $\theta_{1},\dots,\theta_{n}\in\mathsf{Con}_{\mathsf{K}}(\bm{B})$ be such that $\text{id}_{B}=\theta_{1}\cap\dots\cap\theta_{n}$. Let also $\phi_{i}=\theta_{1}\cap\dots\cap\theta_{i-1}\cap\theta_{i+1}\dots\cap\theta_{n}$ for each $i\leqslant n$. We will show that $\phi_{i}=\text{id}_{B}$ for some $i\leqslant n$. Suppose the contrary, with a view to contradiction. Claim 7.11 yields $a_{1},\dots,a_{n}\in A$ such that $\langle a_{i},b\rangle\in\phi_{i}$ for every $i\leqslant n$. By assumption $\mathsf{K}$ has a near unanimity term $t(x_{1},\dots,x_{n})$. We will prove that
@@
In view of Claim 7.12 and Proposition 7.9, there exist $\theta_{1},\dots,\theta_{n-1}\in\mathsf{Irr}_{\mathsf{K}}(\bm{B})$ such that $\text{id}_{B}=\theta_{1}\cap\dots\cap\theta_{n-1}$. Therefore, we can apply Proposition 2.7 obtaining a subdirect embedding $h\colon\bm{B}\to\bm{B}/\theta_{1}\times\dots\times\bm{B}/\theta_{n-1}$. Let $\bm{B}_{i}=\bm{B}/\theta_{i}$ for each $i\leqslant n-1$. By replacing $\bm{A}$ and $\bm{B}$ by their isomorphic images $h[\bm{A}]$ and $h[\bm{B}]$, respectively, we may assume that $\bm{A}\leqslant\bm{B}\leqslant\bm{B}_{1}\times\dots\times\bm{B}_{n-1}$. Notice that each $\bm{B}_{i}=\bm{B}/\theta$ is finitely generated because so is $\bm{B}$. Furthermore, from Proposition 2.10 and $\theta_{i}\in\mathsf{Irr}_{\mathsf{K}}(\bm{B})$ it follows that $\bm{B}_{i}\in\mathsf{K}_{\textsc{rfsi}}$. Lastly, as $b\in\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})-A$ and $\bm{B}\leqslant\bm{B}_{1}\times\dots\times\bm{B}_{n-1}$, Corollary 4.6(ii) allows us to conclude that $b\in\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B}_{1}\times\dots\times\bm{B}_{n-1})-A$. Hence, $\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B}_{1}\times\dots\times\bm{B}_{n-1})\neq A$. ∎
The literature on epimorphisms contains two variants of Theorem 7.3 in which the class $\mathsf{K}$ is required to be an arithmetical variety with the property that the class of its finitely subdirectly irreducible members is closed under ultraproducts and nontrivial subalgebras [23, Thm. 6.8] or only a congruence permutable variety [29, Thm. 5.3]. The first variant deals with the demand that all $\mathsf{K}$-epimorphisms be surjective, while the second with the weaker demand that all $\mathsf{K}$-epimorphisms between finitely generated algebras be surjective called the weak epimorphism surjectivity property. In both cases, the conclusion is that failures of the relevant property are witnessed by counterexamples of the form $\bm{A}\leqslant\bm{B}$ where $\bm{B}$ is a finitely subdirectly irreducible member of $\mathsf{K}$. The possibility of obtaining similar results for the strong epimorphism surjectivity property is prevented by the following example. However, we will show in Corollary 7.16 that, under the amalgamation property, the above mentioned result for congruence permutable varieties becomes available in the context of the strong epimorphism surjectivity property as well.
@@
The next result is well known (see, e.g., [14, Thm. 1.3]). We provide a novel and short proof using the characterization of dominions in the presence of the amalgamation property established in Corollary 4.14.
@@
The implication from left to right is straightforward. So, let us assume that $\mathsf{K}$ has the weak epimorphism surjectivity property. We will show that $\mathsf{K}$ has the strong epimorphism surjectivity property using Proposition 6.6. To this end, consider $\bm{A}\leqslant\bm{B}\in\mathsf{K}$ and $b\in\mathsf{dom}_{\mathsf{K}}(\bm{A},\bm{B})$. Since $\mathsf{K}$ has the amalgamation property, by Corollary 4.14 there exist $f\in\mathsf{imp}_{\textsc{eq}}(\mathsf{K})$ and $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{B}})\cap A^{n}$ such that $f^{\bm{B}}(a_{1},\dots,a_{n})=b$. Let $\bm{A}^{\prime}=\mathsf{Sg}^{\bm{A}}(a_{1},\dots,a_{n})$ and $\bm{B}^{\prime}=\mathsf{Sg}^{\bm{B}}(a_{1},\dots,a_{n},b)$. Since $f$ is defined by a conjunction of equations, $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{B}^{\prime}})\cap(A^{\prime})^{n}$ and $f^{\bm{B}^{\prime}}(a_{1},\dots,a_{n})=b$. So, Corollary 4.14 implies that $b\in\mathsf{d}_{\mathsf{K}}(\bm{A}^{\prime},\bm{B}^{\prime})$. As $\bm{B}^{\prime}=\mathsf{Sg}^{\bm{B}}(a_{1},\dots,a_{n},b)$ and $a_{1},\dots,a_{n},b\in\mathsf{d}_{\mathsf{K}}(\bm{A}^{\prime},\bm{B}^{\prime})$, we obtain $\mathsf{d}_{\mathsf{K}}(\bm{A}^{\prime},\bm{B}^{\prime})=B^{\prime}$. Therefore, the inclusion map $\bm{A}^{\prime}\to\bm{B}^{\prime}$ is an epimorphism. Since $\bm{A}^{\prime},\bm{B}^{\prime}$ are finitely generated members of $\mathsf{K}$ and $\mathsf{K}$ has the weak epimorphism surjectivity property, it follows that $\bm{A}^{\prime}=\bm{B}^{\prime}$, and hence $b\in B^{\prime}=A^{\prime}\subseteq A$. We have shown that $\mathsf{dom}_{\mathsf{K}}(\bm{A},\bm{B})=A$. Thus, $\mathsf{K}$ has the strong epimorphism surjectivity property. ∎
@@
The implication from left to right is straightforward. On the other hand, if every finitely generated $\bm{B}\in\mathsf{K}_{\textsc{fsi}}$ lacks proper $\mathsf{K}$-epic subalgebras, then Theorem 7.15 guarantees that $\mathsf{K}$ has the weak epimorphism surjectivity property, which by Theorem 7.14 implies that $\mathsf{K}$ has the strong epimorphism surjectivity property as well. ∎
@@
Assume that $\mathsf{K}=\mathbb{V}(\mathsf{K}_{1}\cup\dots\cup\mathsf{K}_{n})$, where each $\mathsf{K}_{i}$ is a variety with the weak epimorphism surjectivity property. Suppose, with a view to contradiction, that $\mathsf{K}$ lacks this property. By Theorem 7.15 this implies that there exists a finitely generated $\bm{B}\in\mathsf{K}_{\textsc{fsi}}$ with a finitely generated subalgebra $\bm{A}\leqslant\bm{B}$ that is proper and $\mathsf{K}$-epic. Applying 2.12, we obtain that $\bm{B}\in\mathbb{H}\mathbb{S}\mathbb{P}_{\!\textsc{{u}}}(\mathsf{K}_{1}\cup\dots\cup\mathsf{K}_{n})$. By [11, Thm. 5.6] we have $\mathbb{P}_{\!\textsc{{u}}}(\mathsf{K}_{1}\cup\dots\cup\mathsf{K}_{n})=\mathbb{P}_{\!\textsc{{u}}}(\mathsf{K}_{1})\cup\dots\cup\mathbb{P}_{\!\textsc{{u}}}(\mathsf{K}_{n})$. Therefore, $\bm{B}\in\mathbb{H}\mathbb{S}\mathbb{P}_{\!\textsc{{u}}}(\mathsf{K}_{1})\cup\dots\cup\mathbb{H}\mathbb{S}\mathbb{P}_{\!\textsc{{u}}}(\mathsf{K}_{n})\subseteq\mathsf{K}_{1}\cup\dots\cup\mathsf{K}_{n}$. Then $\bm{B}\in\mathsf{K}_{i}$ for some $i\leqslant n$. But $\mathsf{K}_{i}$ has the weak epimorphism surjectivity property by assumption, whence $\bm{A}\leqslant\bm{B}$ cannot be a $\mathsf{K}_{i}$-epic subalgebra. As $\mathsf{K}_{i}\subseteq\mathsf{K}$, in particular it follows that $\bm{A}\leqslant\bm{B}$ cannot be a $\mathsf{K}$-epic subalgebra either. But this contradicts the assumption and thus completes the proof. ∎